What is the resistance of a copper bus bar 9.15 meter long and 7.62 cm by 1.27

cm in cross section? Use 1.78 μΩ-cm specific resistance for copper.


1
Expert's answer
2021-10-11T16:20:25-0400

"l=9.15m=915cm"

Area, "A=7.62\times 1.27=9.677cm^2"

"\rho=1.78\times 10^{-6}\Omega-cm"

"Resistance, R=\rho \frac{l}{A}=1.78\times 10^{-6} \frac{915}{9.677}=1.683\times 10^{-4} \Omega"